# Admissibility of Stein Shrinkage for Batch Normalization in the Presence of Adversarial Attacks

### v1

@@ Appendix A Technical Appendices and Supplementary Material
Now,
$\mathbb{E}(V)=\mathbb{E}\left(\sqrt[p]{\prod_{i=1}^{p}\hat{\sigma}_{z,i}^{2}}\right)=\mathbb{E}\left(\exp\left(\frac{1}{p}\sum_{i=1}^{p}\log\hat{\sigma}_{z,i}^{2}\right)\right)\geq$
Again, using Jensen’s inequality for $f(x)=\exp{x}$
$\geq\exp\left(\frac{1}{p}\sum_{i=1}^{p}\mathbb{E}(\log\sigma_{z,i}^{2})\right)\geq$
Since $\hat{\sigma}_{y,i}^{2}+W_{i}$ is nonnegative with high probability
$\geq\exp\left(\frac{1}{p}\sum_{i=1}^{p}\mathbb{E}(\log\sigma_{x,i}^{2})\right).$
We know that $\hat{\sigma}^{2}_{x,i}\sim\mathrm{Gamma}(\alpha,\beta_{i})$, so $\mathbb{E}(\log\hat{\sigma}^{2}_{x,i})=\log\beta_{i}+\psi(\alpha),$ where $\psi(\alpha)=\mathbb{E}(\log u)$ is digamma function, $u\sim\mathrm{Gamma}(\alpha,1).$ For $\alpha>0$ we can bound $\psi(\alpha)\geq\log\alpha-\frac{1}{\alpha}.$ Hence,
$\mathbb{E}(\log\sigma_{x,i}^{2})\ \geq\log\beta_{i}+\log\alpha-\frac{1}{\alpha}.$
Plug in this inequality to the derivations above we get:
$\mathbb{E}(V)\geq\exp\left(\frac{1}{p}\sum_{i=1}^{p}\left(\log\beta_{i}+\log\alpha-\frac{1}{\alpha}\right)\right)=\alpha\cdot\exp{\left(-\frac{1}{\alpha}\right)}\left(\prod_{i=1}^{p}\beta_{i}\right)^{\frac{1}{p}}$
For (13) we get:
$\frac{1}{\sqrt{\mathbb{E}(V^{2})}}\leq\frac{1}{\mathbb{E}(V)}\leq\frac{1}{\alpha\cdot\exp{\left(-\frac{1}{\alpha}\right)}\left(\prod_{i=1}^{p}\beta_{i}\right)^{\frac{1}{p}}}$
Also, we can directly calculate numerator of (12):
@@
$=\sqrt{\alpha\sum_{i=1}^{p}\beta_{i}^{2}+\alpha^{2}(\sum_{i=1}^{p}\beta_{i})^{2}}$
Therefor, for (12) we get upper bound:
$\frac{\mathbb{E}\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\cdot V\right)}{\mathbb{E}(V^{2})}\leq\frac{\sqrt{\alpha\sum_{i=1}^{p}\beta_{i}^{2}+\alpha^{2}(\sum_{i=1}^{p}\beta_{i})^{2}}}{\alpha\cdot\exp{\left(-\frac{1}{\alpha}\right)}\left(\prod_{i=1}^{p}\beta_{i}\right)^{\frac{1}{p}}.}$
Using this inequality and (11) finally for (10) we get:
@@
$=\frac{2}{\alpha p+1}\frac{\sqrt{\frac{1}{\alpha}\sum_{i=1}^{p}\beta_{i}^{2}+(\sum_{i=1}^{p}\beta_{i})^{2}}}{\left(\prod_{i=1}^{p}\beta_{i}\right)^{\frac{1}{p}}}\exp{\left(\frac{1}{\alpha}\right)}-\frac{2}{\alpha+1}$
@@
Hence,
@@
$\leq p\frac{\sqrt{\frac{1}{\alpha}\sum_{i=1}^{p}\beta_{i}^{2}+(\sum_{i=1}^{p}\beta_{i})^{2}}}{\sum_{i=1}^{p}\beta_{i}}=p\sqrt{\frac{\frac{1}{\alpha}\sum_{i=1}^{p}\beta_{i}^{2}+(\sum_{i=1}^{p}\beta_{i})^{2}}{(\sum_{i=1}^{p}\beta_{i})^{2}}}<p\sqrt{\frac{1}{\alpha}+1}.$
So, dominance holds for
@@
Note that under regular conditions Stein’s estimate dominance holds when constant $c$ satisfies
@@ Appendix B Experiments
We will, firstly, provide derivations of Lasso and Ridge shrinkage estimators for mean and variance.
Lasso mean shrinkage. We consider the following penalized least squares problem:
$\hat{\mu}=argmin_{\mu}\left(\sum_{i=1}^{n}(x_{i}-\mu)^{2}+\lambda|\mu|\right),$
where $\lambda>0$ is a regularization parameter controlling the amount of shrinkage. This is equivalent to applying an $l_{1}$ – penalty to the mean and the solution is given by the soft-thresholding operator:
$\hat{\mu}=sign(\bar{x})\max{(0,|\bar{x}|-\frac{\lambda}{2n}),}$
where $\bar{x}=\frac{1}{n}\sum_{i=1}^{n}x_{i}$ is the sample mean. This form encourages shrinkage of small mean values towards zero, potentially setting them exactly to zero when $\lambda$ is large.
Lasso variance shrinkage. A similar penalized formulation can be used for estimating the variance, leading to a soft-thresholding-type estimator for sample variance $s^{2}$:
$\hat{\sigma}^{2}=\max{(0,s^{2}-\frac{\lambda}{2}),}$
where $\lambda>0$ is regularization hyperparameter, $s^{2}=\frac{1}{n}\sum_{i}(x_{i}-\bar{x})^{2}.$
Ridge mean shrinkage. We consider the following penalized least squares problem:
$\hat{\mu}=argmin_{\mu}\left(\sum_{i=1}^{n}(x_{i}-\mu)^{2}+\lambda\mu^{2}\right),$
where $\lambda>0$ is a regularization parameter controlling the amount of shrinkage. This is equivalent to applying an $l_{2}$ – penalty to the mean. Taking the derivative and setting it to zero yields the following solution:
$\hat{\mu}=\frac{\sum_{i}x_{i}}{n+\lambda}.$
Ridge variance shrinkage. Similarly, we can obtain the ridge estimate for the variance:
$\hat{\sigma}^{2}=\frac{s^{2}}{1+\lambda},$
where $s^{2}$ is the sample variance and $\lambda>0$ is the regularization parameter.
Now let us discuss reproducibility of experiments on each dataset.
B.1 CIFAR10 dataset
We conducted experiments on the CIFAR-10 dataset, which consists of 50,000 training and 10,000 test images across 10 classes. We simulate sub-gaussian noise by perturbing the input features with zero-mean noise of varying magnitudes $(0-30\%)$, applied independently to each image during validation and testing. Note that in this appendix, we present an additional set of results not in the main body of the paper namely, the results of applying Stein’s shrinkage estimation to just the mean parameter (no variance shrinkage) used in the BN and included in the Table 4.
We used the standard ResNet-9 architecture, modifying only the batch normalization layers according to the method under consideration. For data splits, we used the standard predefined training and test sets. Training time was approximately $10$ minutes per run, with negligible variation across methods.
All experiments were conducted using a single NVIDIA A100 GPU, $4$ CPUs, and $16$ GB of RAM. Each experiment was repeated $11$ times to estimate the standard deviation of the performance metrics, and we tested three different batch sizes $(32,64,128)$, resulting in a total training time of approximately $1800$ minutes.
We used SGD with Nesterov acceleration and default parameters, except for the learning rate, which was tuned separately for each method within the range $(10^{-5},10^{-2})$. We employed early stopping based on validation accuracy on data without noise to prevent overfitting. We used cross-entropy loss for training, and report classification accuracy as the main evaluation metric. Finally, for the Lasso and Ridge shrinkage methods, we selected the regularization parameter $\lambda\in(10^{-3},10^{-1})$ yielding the best performance on validation dataset without noise.
Method 0% noise 3% noise 10% noise 15% noise 20% noise 30% noise
Standard BN (32) 79.2 $\pm$ 0.28 29.83 $\pm$ 2.51 17.19 $\pm$ 1.6 15.28 $\pm$ 1.38 14.38 $\pm$ 0.83 14.22 $\pm$ 1.32
Standard BN (64) 81.7 $\pm$ 0.27 30.17 $\pm$ 2.73 19.19 $\pm$ 2.05 17.10 $\pm$ 1.54 15.54 $\pm$ 2.07 13.7 $\pm$ 0.93
Standard BN (128) 79.8 $\pm$ 0.3 29.05 $\pm$ 3.63 19.48 $\pm$ 2.11 16.23 $\pm$ 1.36 19.05 $\pm$ 3.63 19.48 $\pm$ 2.11
Our BN (32) 80.89 $\pm$ 0.51 60.7 $\pm$ 0.96 44.55 $\pm$ 1.4 35.34 $\pm$ 1.6 26.67 $\pm$ 3.38 25.66 $\pm$ 3.63
Our BN (64) 80.78 $\pm$ 0.54 61.7 $\pm$ 0.81 45.17 $\pm$ 1.43 36.19 $\pm$ 0.57 32.05 $\pm$ 0.58 25.26 $\pm$ 0.69
Our BN (128) 79.1 $\pm$ 0.3 60.49 $\pm$ 1.06 44.96 $\pm$ 1.3 36.27 $\pm$ 1.01 29.2 $\pm$ 4.67 28.12 $\pm$ 3.32
Mean corrected BN (32) 78.7 $\pm$ 1.32 56.79 $\pm$ 0.87 40.13 $\pm$ 1.19 31.57 $\pm$ 1.19 25.82 $\pm$ 3.65 24.77 $\pm$ 3.97
Mean corrected BN (64) 80.6 $\pm$ 0.4 54.15 $\pm$ 1.26 38.24 $\pm$ 0.66 30.17 $\pm$ 1.37 26.78 $\pm$ 1.42 20.85 $\pm$ 1.13
Mean corrected BN (128) 79.9 $\pm$ 0.48 49.2 $\pm$ 1.35 32.96 $\pm$ 0.99 25.97 $\pm$ 1.2 25.82 $\pm$ 3.65 24.77 $\pm$ 3.97
Khoshsirat et al. BN [15] (32) 76.82 $\pm$ 1.01 57. 49 $\pm$ 1.71 41.93 $\pm$ 1.46 33.20 $\pm$ 1.73 27.10 $\pm$ 3.5 25.39 $\pm$ 3.72
Khoshsirat et al. BN [15] (64) 79.84 $\pm$ 1.18 58.71 $\pm$ 1.48 43.75 $\pm$ 1.67 33.92 $\pm$ 2.4 30.75 $\pm$ 1.32 23.36 $\pm$ 1.20
Khoshsirat et al. BN [15] (128) 79.12 $\pm$ 0.27 60.02 $\pm$ 1.11 44.85 $\pm$ 1.47 36.35 $\pm$ 1.47 29.15 $\pm$ 4.25 28.34 $\pm$ 3.26
Lasso corrected BN (32) 78.4 $\pm$ 1.0 27.50 $\pm$ 3.88 16.43 $\pm$ 2.42 14.79 $\pm$ 2.08 14.9 $\pm$ 1.59 14.28 $\pm$ 1.32
Lasso corrected BN (64) 80.3 $\pm$ 0.6 28.42 $\pm$ 3.75 17.54 $\pm$ 3.29 16.47 $\pm$ 1.44 16.27 $\pm$ 0.9 13.91 $\pm$ 1.12
Lasso corrected BN (128) 79.3 $\pm$ 0.38 25.79 $\pm$ 6.26 17.77 $\pm$ 3.04 15.22 $\pm$ 1.45 14.27 $\pm$ 1.91 13.96 $\pm$ 1.93
Ridge corrected BN(32) 78.4 $\pm$ 0.69 30.13 $\pm$ 4.58 17.49 $\pm$ 2.89 15.29 $\pm$ 2.18 14.68 $\pm$ 2.13 14.37 $\pm$ 0.69
Ridge corrected BN(64) 80.6 $\pm$ 0.41 27.36 $\pm$ 3.8 18.32 $\pm$ 3.09 15.0 $\pm$ 1.22 15.62 $\pm$ 0.56 13.4 $\pm$ 0.61
Ridge corrected BN(128) 79.9 $\pm$ 0.48 28.36 $\pm$ 3.24 19.53 $\pm$ 1.82 16.67 $\pm$ 1.4 13.48 $\pm$ 1.64 13.77 $\pm$ 2.18
Table 4: CIFAR-10 Test data classification performance using BN methods under varying subgaussian noise and batch sizes respectively. Values depict mean accuracy $\pm$ standard deviation.
B.2 Cityscapes dataset
We performed segmentaion experiments on Cityscapes dataset – high-resolution urban scene segmentation with fine-grained object boundaries. The dataset contains 5,000 finely annotated images from 50 cities with 19 semantic classes used for evaluation. In all our experiments we injected subgaussian noise into the feature maps during validation and testing to simulate corrupted statistics. Note that in this appendix, we present an additional set of results not in the main body of the paper namely, the results of applying Stein’s shrinkage estimation to just the mean parameter (no variance shrinkage) used in the BN and included in the Table 5. We use the HRNetV2-W16 architecture (where 16 denotes the width of the high-resolution convolution) as our baseline model. The original training set was randomly split into $80\%$ for training and $20\%$ for validation, while the original validation set was used as the test set. All images were resized from $1024\times 2048$ to $512\times 1024$ for computational efficiency. To account for variability and estimate the standard deviation of our results, each experiment was repeated $7$ times for each batch size $(16,32).$ Each run took approximately $1.5$ hours on a single NVIDIA A100 GPU with $64$ GB of RAM and $8$ CPU cores, resulting in a total training time of around 126 hours.
@@ B.2 Cityscapes dataset
We trained the models using stochastic gradient descent (SGD) with Nesterov acceleration and default settings, except for the learning rate, which was tuned separately for each method within the range $(10^{-5},10^{-2})$. Early stopping based on the mean Intersection over Union (mIoU) score on validation data without noise was employed to prevent overfitting. We used cross-entropy loss for training and report mIoU as the primary evaluation metric. Finally, for the Lasso and Ridge shrinkage methods, the regularization parameter $\lambda\in(10^{-3},10^{-1})$ was selected based on performance on the validation set without noise.
Method 0% noise 3% noise 10% noise 15% noise 20% noise 30% noise
Standard BN (8) 68.1 $\pm$ 0.9 8.6 $\pm$ 0.8 7 $\pm$ 0.9 6.14 $\pm$ 1.6 5.7 $\pm$ 2.0 4.8 $\pm$ 2.9
Standard BN (16) 67.4 $\pm$ 0.9 9.9 $\pm$ 1.7 7.6 $\pm$ 2.1 6.8 $\pm$ 2.4 6.4 $\pm$ 2.5 5.7 $\pm$ 2.6
Our BN (8) 68.8 $\pm$ 0.1 39.6 $\pm$ 2.3 32.5 $\pm$ 2.6 29.9 $\pm$ 2.4 28.9 $\pm$ 2.3 27.2 $\pm$ 1.8
Our BN (16) 68.4 $\pm$ 0.4 36.3 $\pm$ 2.7 29.4 $\pm$ 1.5 27.1 $\pm$ 1.5 26.3 $\pm$ 1.7 25.1 $\pm$ 2.2
Mean corrected BN (8) 67.80 $\pm$ 0.2 39.13 $\pm$ 2.6 31.8 $\pm$ 2.1 29.1 $\pm$ 2 28.1 $\pm$ 2 26.4 $\pm$ 2.2
Mean corrected BN (16) 68.1 $\pm$ 0.5 39.07 $\pm$ 1.7 32.8 $\pm$ 1 30.8 $\pm$ 1.2 30 $\pm$ 1.3 28.6 $\pm$ 1.2
Khoshsirat et al. BN [15] (8) 68.4 $\pm$ 0.3 36.4 $\pm$ 2.8 28.5 $\pm$ 3.5 25.9 $\pm$ 3.5 24.8 $\pm$ 3.2 23.3 $\pm$ 3
Khoshsirat et al. BN [15] (16) 67.1 $\pm$ 0.3 36.2 $\pm$ 0.7 28.7 $\pm$ 1.3 26.3 $\pm$ 1.3 25.5 $\pm$ 1.2 24.4 $\pm$ 1.2
Lasso corrected BN (8) 65.9 $\pm$ 0.2 9.7 $\pm$ 4.2 6.2 $\pm$ 2.5 5.1 $\pm$ 2.2 4.6 $\pm$ 2.1 3.9 $\pm$ 1.9
Lasso corrected BN (16) 64.9 $\pm$ 0.9 8 $\pm$ 4.3 5.5 $\pm$ 3.2 4.7 $\pm$ 3.1 4.3 $\pm$ 3.1 3.7 $\pm$ 3
Ridge corrected BN(8) 68.4 $\pm$ 0.9 8.4 $\pm$ 1.4 7.3 $\pm$ 1.6 7.1 $\pm$ 1.9 7 $\pm$ 2.1 6 $\pm$ 2.3
Ridge corrected BN(16) 66.8 $\pm$ 0.1 6.3 $\pm$ 1.9 4.6 $\pm$ 2.1 4.2 $\pm$ 2.5 4.1 $\pm$ 2.6 3.9 $\pm$ 2.7
Table 5: Performance of HRNetV2 with different BN methods under varying subgaussian noise and batch sizes respectively. Values depict mean of mIoU scores $\pm$ standard deviation.
B.3 Details of Application to PPMI Data
In this experiment, the goal is to demonstrate the performance of our BN with Stein shrinkage on 3D diffusion tensor magnetic resonance images (DT-MRI or in short DTI) acquired from Parkinson’s Disease (PD) affected brains and normal brains (Controls). This data was obtained from the PPMI (Parkinson’s Progressive Markers Initiative) portal after a formal request was filled out. Please see data access citation and acknowledgment below for this data.
DTI is a non-invasive MRI technique that allows one to infer neuronal connectivity patterns in the tissue being imaged. The PPMI data was subjected to a standard pre-processing pipeline involving Eddy currents and motion correction etc., using the FSL software [13]. DTIs are matrix-valued images in which each voxel takes on a value that is a $(3,3)$ symmetric positive definite (SPD) matrix representing the covariance matrix of a zero-mean local Gaussian at the voxel. These DTIs are further processed to extract scalar-valued feature maps, called fractional anisotropy (FA) maps. The FA is a scalar function of the three eigen values of the diffusion tensor at each voxel in the DTI. Harmonization is applied to these scalar feature maps to remove site effects caused by distinct acquisition protocols and equipment used. Harmonization was achieved using the method called ComBat adapted to diffusion tensor images and associated software presented in [12]. It should be noted that there are other sophisticated hamronization methods that one can apply to the data but this one is simple. The data set is imbalanced and consists of $490$ (130 PD and 360 Controls) FA maps. To address this, we use mixup-based[31] augmentation with sampling bias favoring the minority class (PD), this approach allowed us to generate additional training examples. We follow the same procedure in our other experiments for the setup and use a 5-fold cross validation to compare 3D CNN models with different variants of 3D BatchNorm layer. The architecture consists of four 3D convolutional layers, each followed by a 3D BatchNorm layer, and three linear layers for binary classification. The 3D input tensors have spatial dimensions of $(128\times 128)$ and a depth of $64$. We conduct experiments using two batch sizes ($10$ and $16$) and optimize the binary cross-entropy loss. Note that in this appendix, we present an additional set of results not in the main body of the paper namely, the results of applying Stein’s shrinkage estimation to just the mean parameter (no variance shrinkage) used in the BN and included in the Table 6. We trained the models using stochastic gradient descent (SGD) with Nesterov acceleration and default settings, except for the learning rate, which was tuned separately for each method within the range $(10^{-5},10^{-2})$. Early stopping based on the accuracy score on validation data without noise was employed to prevent overfitting. Finally, for the Lasso and Ridge shrinkage methods, the regularization parameter $\lambda\in(10^{-3},10^{-1})$ was selected based on performance on the validation set without noise. All experiments were conducted using a single NVIDIA A100 GPU, $4$ CPUs, and $16$ GB of RAM. Each configuration take around $20$ minutes, which end up to the 20 hours for the whole experiments.
Method 3% noise 10% noise 15% noise 20% noise 30% noise
Standard BN (10) 71.2 $\pm$ 1.6 64.7 $\pm$ 4.3 62.3 $\pm$ 5 61.3 $\pm$ 3.9 60.7 $\pm$ 8.2
Standard BN (16) 72.6 $\pm$ 1.5 63.7 $\pm$ 6.7 63.5 $\pm$ 4.6 62.3 $\pm$ 5.6 59.1 $\pm$ 5.4
Our BN (10) 71.8 $\pm$ 2.8 68 $\pm$ 3.6 68 $\pm$ 3.9 66.8 $\pm$ 5.2 66.5 $\pm$ 6.9
Our BN (16) 72.4 $\pm$ 2.8 65.7 $\pm$ 3.6 63.5 $\pm$ 6 61.5 $\pm$ 5.3 60.9 $\pm$ 7.2
Mean corrected BN (10) 68.6 $\pm$ 7.3 64.7 $\pm$ 6.1 64.1 $\pm$ 5.7 63.7 $\pm$ 6.1 62.3 $\pm$ 4.3
Mean corrected BN (16) 69.2 $\pm$ 7.9 62.9 $\pm$ 12.2 61.1 $\pm$ 7.5 59.7 $\pm$ 10.9 58.7 $\pm$ 8.2
Khoshsirat et al. BN [15] (10) 71.4 $\pm$ 2.4 67.6 $\pm$ 4.5 66.4 $\pm$ 4.5 65.3 $\pm$ 7 64.9 $\pm$ 7.2
Khoshsirat et al. BN [15] (16) 70.8 $\pm$ 1.9 67.2 $\pm$ 3.9 65.1 $\pm$ 6.1 64.3 $\pm$ 3.7 63.5 $\pm$ 4.6
Lasso corrected BN (10) 64.9 $\pm$ 6.9 62.5 $\pm$ 9.4 62.3 $\pm$ 12.9 60.7 $\pm$ 9.3 57.5 $\pm$ 7.1
Lasso corrected BN (16) 65.3 $\pm$ 3.9 64.5 $\pm$ 5.2 62.9 $\pm$ 4.9 59.1 $\pm$ 4.6 57.2 $\pm$ 4.4
Ridge corrected BN(10) 68.8 $\pm$ 3.6 64.7 $\pm$ 4.1 64.5 $\pm$ 6.5 62.9 $\pm$ 4 61.7 $\pm$ 5.3
Ridge corrected BN(16) 71 $\pm$ 1.8 64.3 $\pm$ 7.4 63.7 $\pm$ 3.3 62.1 $\pm$ 5.5 61.1 $\pm$ 2.6
Table 6: 3D CNN performance with different BN methods under varying subgaussian noise and batch sizes respectively on FA maps data. Values depict mean of accuracies $\pm$ standard deviation.

### v2

@@ Appendix A Technical Appendices and Supplementary Material
Again, applying Jensen’s inequality to the function $f(x)=\exp{x}$ in the following expression, we get,
$\mathbb{E}(V)=\mathbb{E}\left(\sqrt[p]{\prod_{i=1}^{p}\hat{\sigma}_{z,i}^{2}}\right)=\mathbb{E}\left(\exp\left(\frac{1}{p}\sum_{i=1}^{p}\log\hat{\sigma}_{z,i}^{2}\right)\right)\geq\exp\left(\frac{1}{p}\sum_{i=1}^{p}\mathbb{E}(\log\sigma_{z,i}^{2})\right)\geq$
Since $\hat{\sigma}_{y,i}^{2}+W_{i}$ is nonnegative with high probability (see lemma 6) we get:
$\mathbb{E}(V)\geq\exp\left(\frac{1}{p}\sum_{i=1}^{p}\mathbb{E}(\log\sigma_{x,i}^{2})\right).$ (32)
Recall that $\hat{\sigma}^{2}_{x,i}\sim\mathrm{Gamma}(\alpha,\beta_{i})$, so $\mathbb{E}(\log\hat{\sigma}^{2}_{x,i})=\log\beta_{i}+\psi(\alpha),$ where $\psi(\alpha)=\mathbb{E}(\log u)$ is a digamma function, $u\sim\mathrm{Gamma}(\alpha,1).$ For $\alpha>0$ we can use the bound $\psi(\alpha)\geq\log\alpha-\frac{1}{\alpha}.$ Hence,
$\mathbb{E}(\log\sigma_{x,i}^{2})\geq\log\beta_{i}+\log\alpha-\frac{1}{\alpha}.$ (33)
Substituting the bound (33) into (32) we obtain the following inequality:
$\mathbb{E}(V)\geq\exp\left(\frac{1}{p}\sum_{i=1}^{p}\left(\log\beta_{i}+\log\alpha-\frac{1}{\alpha}\right)\right)=\alpha\cdot\exp{\left(-\frac{1}{\alpha}\right)}\left(\prod_{i=1}^{p}\beta_{i}\right)^{\frac{1}{p}}$ (34)
Now, using (34) for (31) we get:
$\frac{1}{\sqrt{\mathbb{E}(V^{2})}}\leq\frac{1}{\mathbb{E}(V)}\leq\frac{1}{\alpha\cdot\exp{\left(-\frac{1}{\alpha}\right)}\left(\prod_{i=1}^{p}\beta_{i}\right)^{\frac{1}{p}}}$ (35)
Also, we can directly calculate the numerator of (30):
@@
$\sqrt{\alpha\sum_{i=1}^{p}\beta_{i}^{2}+\alpha^{2}(\sum_{i=1}^{p}\beta_{i})^{2}}$ (36)
Therefore, combining (36) and (35) we get an upper bound for (30):
$\frac{\mathbb{E}\left(\sum_{i=1}^{p}\hat{\sigma}_{x,i}^{2}\cdot V\right)}{\mathbb{E}(V^{2})}\leq\frac{\sqrt{\alpha\sum_{i=1}^{p}\beta_{i}^{2}+\alpha^{2}(\sum_{i=1}^{p}\beta_{i})^{2}}}{\alpha\cdot\exp{\left(-\frac{1}{\alpha}\right)}\left(\prod_{i=1}^{p}\beta_{i}\right)^{\frac{1}{p}}.}$ (37)
Finally, for (27) using inequality (37) and (28) we get:
@@
$\frac{2}{\alpha p+1}\frac{\sqrt{\frac{1}{\alpha}\sum_{i=1}^{p}\beta_{i}^{2}+(\sum_{i=1}^{p}\beta_{i})^{2}}}{\left(\prod_{i=1}^{p}\beta_{i}\right)^{\frac{1}{p}}}\exp{\left(\frac{1}{\alpha}\right)}-\frac{2}{\alpha+1}$ (38)
@@
Hence, we can write a bound for the first term in (38):
@@
$\leq p\frac{\sqrt{\frac{1}{\alpha}\sum_{i=1}^{p}\beta_{i}^{2}+(\sum_{i=1}^{p}\beta_{i})^{2}}}{\sum_{i=1}^{p}\beta_{i}}=p\sqrt{\frac{\frac{1}{\alpha}\sum_{i=1}^{p}\beta_{i}^{2}+(\sum_{i=1}^{p}\beta_{i})^{2}}{(\sum_{i=1}^{p}\beta_{i})^{2}}}<p\sqrt{\frac{1}{\alpha}+1}.$ (39)
According to (38) in combination with (39) dominance holds for
@@
Note that under regular conditions the dominance of Stein’s estimate holds when constant $c$ satisfies
@@ Appendix B Experiments
Unless otherwise specified, all models were trained with stochastic gradient descent (SGD) with Nesterov acceleration, cross-entropy loss, and early stopping based on validation performance on noise-free data. The learning rate was adjusted separately for each method in the range $(10^{-5},10^{-2})$. For the Lasso and Ridge shrinkage methods, the regularization parameter $\lambda\in(10^{-3},10^{-1})$ was selected based on the noise-free validation performance. All experiments were performed on NVIDIA A100 GPUs with CPU and RAM resources adjusted to the size of the data set. To assess variability, each setting was repeated multiple times and we report averages with standard deviations.
B.1 CIFAR-10
[…]
B.2 Cityscapes
We used the HRNetV2-W16 architecture with images resized from $1024\times 2048$ to $512\times 1024$. The dataset contains 5,000 finely annotated images across 19 classes. We split the original training set into $80\%$ training and $20\%$ validation, and used the original validation set as test data. Sub-Gaussian noise was injected into feature maps during validation and testing. We tested batch sizes $(16,32)$, with each experiment repeated $7$ times. Each run took $\sim$1.5 hours, totaling $\sim$126 GPU hours.
B.3 PPMI Data
We employed 3D CNN architecture consisting of four convolutional layers, each followed by a 3D BN variant under consideration, and three fully connected layers for binary classification. The input tensors had spatial dimensions of $(128\times 128)$ and a depth of $64$. We tested batch sizes of $10$ and $16$, using 5-fold cross-validation. Each configuration took $\sim$20 minutes, for a total of $\sim$20 GPU hours.
B.4 Comparison with data dependent attacks
For data dependent adversarial attack experiments, we use trained models with their best-performing hyperparameters and apply attacks only to the test data. The overall computational setup, including runtime, memory requirements, and hardware, remains the same as in the previous experiments. As mentioned earlier, both FGSM and PGD attacks are evaluated at different perturbation levels whose values range over, ($\epsilon=0.03,0.1,0.15,0.2,0.3$). For PGD, we additionally specify attack hyperparameters: on the CIFAR-10 dataset, we use 5 iterations with a fixed step size $\alpha=1/255$, while on the Cityscapes dataset we use 5 iterations with an adaptive step size $\alpha=\epsilon/T$, where $T$ is the total number of iterations. The corresponding results are reported in Tables 4, 5, 6, and 7 respectively.